OpenStudy (anonymous):

Can someone please help me with this I have no idea how to do this Solve: 3(x + 4) x - 2

OpenStudy (blurbendy):

distribute the 3, then combine like terms

OpenStudy (anonymous):

what do you mean?

OpenStudy (blurbendy):

does the question look like this: 3(x+4)(x-2) ?

OpenStudy (blurbendy):

or like: 3(x+4) + x -2 ?

OpenStudy (anonymous):

it says 3(x+4)(ge) x-2

OpenStudy (blurbendy):

(ge) ???

OpenStudy (anonymous):

3(x+4)\[(\ge) \] x-2

OpenStudy (anonymous):

it wont let me type it all together

OpenStudy (blurbendy):

3(x+4) >= x -2 3x + 12 >= x -2 That's what I mean by distribute. Can you solve from here?

OpenStudy (anonymous):

do i add 2 to both sides?

OpenStudy (blurbendy):

you can. i like to get the variables on one side and the constants on the other. that way we can solve for x by itself.

OpenStudy (anonymous):

okay once i get 14 what do i do from there?

OpenStudy (blurbendy):

I would have solved it like this: 3x + 12 >= x -2 subtract x from both sides 2x + 12 >= -2 subtract 12 from both sides 2x >= -14 divide both sides by 2 x >= -7

OpenStudy (anonymous):

so did i do it backwards?

OpenStudy (blurbendy):

it doesnt matter which way you do it, as long as you get x >= -7. that can also be written as -7 <= x

OpenStudy (anonymous):

oh okay im not understanding what did you do to the 3? but thank you soooo much!

OpenStudy (blurbendy):

3x + 12 >= x - 2 Here I subtracted x from both sides If you do that you end up with: 2x + 12 >= -2 3x - x = 2x

OpenStudy (blurbendy):

do you see?

OpenStudy (anonymous):

so what do you do with the -2?

OpenStudy (blurbendy):

i left it where it was, but i subtracted -12 from both sides so i got: 2x >= -14 -12 - 2 = -14

OpenStudy (anonymous):

ohhh okay i think i get it! thanks so much

OpenStudy (blurbendy):

cool, anytime!